OpenStudy (anonymous):

A ladder id leaning against a building the distance from the bottom of the ladder to the building is 2ft shorter than the length of the ladder. How high up the side of the building is the building is the top of the ladder if the distance is 1ft less than the length of ladder?

OpenStudy (anonymous):

: [a^2 + b^2 = c^2] : (l-1)^2 + (l-2)^2 = l^2 : (l^2-2l+1) + (l^2-4l+4) = l^2 : 2l^2 - 6l + 5 = l^2 : l^2 - 6l + 5 = 0 Then you factorise : (l-5)(l-1) = 0 Therefore, l = 5 or l = 1. : 2l^2

OpenStudy (anonymous):

ladder= 2ft + 1ft ={3ft] ground= 2ft wall= ? Pythagorean Theorem: a^2 + b^2 =c^2 ---------------------------------- 2^2 + b^2 =3^2 4 + b^2 =9 4(-4) + b^2 =9(-4) b^2 =5 b= (square root of:) 5 b= 2.236067977 :: 2.24ft

OpenStudy (anonymous):

So whats the answer?